It has been estimated that a man's heart is about 33 cm below his brain. Considering the density of human blood to be 1.1x10³ kg/m³, what pressure must the heart produce to supply blood to the brain?


1
Expert's answer
2021-05-05T07:41:34-0400

Pressure within the body is an enclosed system and can be calculated by the formula as follows,


Pressure= "h\rho g"

where h= height of fluid column to be pumped

"\rho =" the density of fluid

"g=" the gravitational acceleration assumed as 10 "m\/s^2" ("N\/Kg" )

Pressure = "0.33 \times 1.1 \times 10^3 \times 10"


P= 36330 "N\/m^2"